OpenStudy (anonymous):

I Will Give Medal. Part A: Write the expression x^2 + 7x + 12 as a product of two linear expressions. Show your work and justify each step. Part B: Rewrite x^2 + 4x + 4 as a square of a linear expression. Part C: Do the expressions in parts A and B have a common factor? Justify your answer.

OpenStudy (anonymous):

since \(4\times 3=12\) and \(4+3=7\) the first one factors as \[(x+3)(x+4)\]

OpenStudy (anonymous):

\[x^2+4x+4\] is a perfect square it is the square of \(x+2\) which is another way of saying \[x^2+4x+4=(x+2)^2\]

OpenStudy (anonymous):

doesn't look like they have any common factors to me, since \[(x+3)(x+4)\] and \[(x+2)(x+2)\] have no factors that are the same